A survey found that people keep their televisions an average of 5.2 years. The standard deviation is 0.8 year. Assume the random variable is normally distributed. If you buy a TV for your dorm room and keep it for 3.8 years of school then sell it, what percentile does that put you in?
Answer by Boreal(15235)

z=(sample -µ)/sd
z=(3.8-5.2)/0.8 ; This = -1.4/0.8 = - 1.571
I want z value to the left of -1.571
That is 0.058 or the 6th percentile.